### v1

# Manga109Dialog: A Large-scale Dialogue Dataset for Comics Speaker Detection

Abstract
The expanding market for e-comics has spurred interest in the development of automated methods to analyze comics. For further understanding of comics, an automated approach is needed to link text in comics to characters speaking the words. Comics speaker detection research has practical applications, such as automatic character assignment for audiobooks, automatic translation according to characters’ personalities, and inference of character relationships and stories. To deal with the problem of insufficient speaker-to-text annotations, we created a new annotation dataset Manga109Dialog [^1] Dataset and code are available at https://github.com/manga109/public-annotations. based on Manga109. Manga109Dialog is the world’s largest comics speaker annotation dataset, containing 132,692 speaker-to-text pairs. We further divided our dataset into different levels by prediction difficulties to evaluate speaker detection methods more appropriately. Unlike existing methods mainly based on distances, we propose a deep learning-based method using scene graph generation models. Due to the unique features of comics, we enhance the performance of our proposed model by considering the frame reading order. We conducted experiments using Manga109Dialog and other datasets. Experimental results demonstrate that our scene-graph-based approach outperforms existing methods, achieving a prediction accuracy of over 75%.
1 Introduction
The market for e-comics has been expanding rapidly with the popularization of digital devices. In 2022, e-comics made up 66.2% of the Japanese comics market shuppankagaku2023 . The development of e-comics has led to increased interest in automated methods to analyses comics. To support these techniques, reliable large-scale dialogue datasets are required for improved computational understanding of comics. Additionally, automated methods to detect characters to whom text is attributed are necessary for effective speaker detection.
Figure 1: An example of speaker detection. ©Akamatsu Ken.
An example of speaker detection is shown in Figure 1. Given an image from comics, the system can detect characters, text, and frame regions, and can then automatically predict the character to which the text is attributed. This research can be applied to various tasks, such as automatic character assignment for text-to-speech reading, automatic translation according to characters’ personalities, inference of character relationships and stories, and automated generation of scenarios.
Due to the lack of dialogue datasets and some limitations of existing annotations, we constructed more accurate and rich annotations, Manga109Dialog. To the best of our knowledge, this is the largest comics dialogue dataset ever created. We divided the test set into two subsets based on the prediction difficulty.
To evaluate the reliability and accuracy of Manga109Dialog, we conducted speaker detection experiments on it. Previous approaches were primarily based on the rule that the character nearest to a given text is most likely to be the speaker rigaud2015speech ; abe2020 . However, due to the unique and complex structure of comics, they may have difficulty making correct predictions in some complex cases. As an example, consider the text at bottom right in Figure 1. If we chose the character nearest to the text as the speaker, the boy would be incorrectly predicted as the speaker rather than the girl. Therefore, the relationship between characters and text should be considered to make a correct prediction.
To address this challenge, we propose an innovative deep learning-based approach. We leverage scene graphs, which are widely recognized as one of the most popular methods for describing visual relationships. Scene graph generation (SGG) is the task of detecting objects and their relationships in an image. Since SGG models have performed well on real-world datasets, such as the Visual Genome dataset krishna2017visual , we propose utilizing SGG models for the task of comics speaker detection. Along with the general framework of SGG models, we introduce frame information because we find that the speaker tends to appear in the same frame as the text.
Traditional metrics used to evaluate SGG lack suitability because the amount of speaker-to-text pairs varies from page to page. Besides, prediction results using traditional metrics cannot be compared to those of rule-based methods. Therefore, we present an evaluation metric well suited for this specific task, and demonstrate its superiority over conventional approaches in experimental results on Manga109Dialog. The contributions are summarized as follows.
  • We constructed an annotation dataset of associations between speakers and texts. This is the largest comics dialogue dataset in the world, containing over 130,000 speaker-to-text pairs.
  • We propose a deep learning-based approach for comics speaker detection using SGG models. We enhance the results by introducing frame information in the relationship prediction stage.
  • We introduce a new evaluation metric and a new standard for data division, that is, dividing the dataset according to prediction difficulty.
Experimental results showed that Manga109Dialog provides a challenging yet realistic benchmark for comics speaker detection, and that our proposed approach exhibited 5% better performance compared to conventional rule-based methods. This research aimed to show to what extent speaker detection can be achieved solely with visual information, providing a benchmark for incorporating NLP in future research.
2 Related Work
2.1 Comics Speaker Dataset
Among the many comics available in different forms, only a few can be used for academic research due to copyright issues. eBDtheque guerin2013ebdtheque is a dataset of comics available for direct use. It comprises 100 images from French, American, and Japanese comics, containing 1,550 characters, 1,092 speech balloons, and 887 speaker-to-text pairs. Moreover, a large-scale manga dataset called Manga109 aizawa2020building contains 109 Japanese comics, including 21,142 images, 2,979 characters, and 147,918 texts. Given the lack of available comics datasets associated with speakers and comics texts, Abe et al. abe2020 constructed a dataset based on Manga109 with a total of 121,364 annotated speaker-to-text pairs.
2.2 Comics Speaker Detection
Rule-Based Methods:
Rigaud et al. rigaud2015speech proposed early rule-based speaker prediction method using distance calculation between speech balloon centroids and character centroids. The character with the shortest distance was assumed to be the speaker. They performed experiments on eBDtheque and obtained an accuracy of 78.58%. In the study by Abe et al. on Manga109, four types of information were used abe2020 , including the text-character distance, the characters in the same frame with the text, the direction of the tail of the speech balloon, and language style (such as first-person utterances). Their approach experimentally obtained an accuracy of 70.7%.
Deep Learning-Based Methods:
In recent years, deep learning-based methods have also been developed. Yamamoto et al. yamamoto2018 used features extracted from the input images and metadata to calculate scores for each text-character pair, while other studies explored time-series learning nakagawa2019 and Natural Language Processing yamaguchi2021 . The results of these studies on Manga109 all demonstrate the potential of learning-based methods in speaker prediction. However, their results were only based on a part of the dataset and lacked implementation details. Therefore, a large dataset is needed to evaluate these models under a unified standard.
2.3 Scene Graph Generation (SGG)
Scene graphs were introduced by Johnson et al. johnson2015image as a data structure for describing objects in a scene. To generate a scene graph, an SGG model first detects object regions and their categories, then identifies relationships between objects. Since scene graphs provide rich semantic features, they can distinguish images or videos more accurately and describe them more precisely. Scene graphs have demonstrated success in various visual tasks, including image retrieval, image captioning, and visual question answering johnson2015image ; yao2018exploring ; tang2019learning .
3 Manga109Dialog
Abe et al. abe2020 first created a comics speaker dataset based on the entire Manga109 dataset. Although their study was pioneering and impressive, this dataset still involves some notable limitations. First, they linked the name of the character to the text instead of the bounding box of the speaker. Therefore, we cannot specify a specific bounding box if the same character appears multiple times on a single page. Besides, since annotators prefer making annotations of simple relationships, some more complicated cases are not included in this dataset.
Aiming to produce a more complete and applicable dataset for speaker detection in comics, we constructed a new dataset, Manga109Dialog. Instead of connecting the character name to the text, we connect the bounding box of the speaker to the text. The visual differences between Abe’s dataset and Manga109Dialog are shown as Figure 2.
3.1 Dataset Construction
In this section, we show the details of how we constructed our dataset. Our annotations were mainly based on the following rules.
  • If the character to whom the text is attributed appears only one on the page, link the bounding box of the character to the text regardless of the position.
  • If more than one character is on the page, the bounding box of the speaker in the same frame as the text is linked in priority. If the speaker is not in the same frame, it is determined by the reading order (from top to bottom, right to left). The texts in the last frame of the page, however, are linked to the character in the second-to-last frame.
  • When there is more than one speaker, we link the text to all speakers.
  • Texts spoken by “Others” are annotated if we can specify the bounding box of the speaker, and are excluded otherwise.
Although we make annotations on Manga109, not all texts are our annotation targets. For example, titles and descriptions were not annotated. Texts designed as “Narration” or with unknown speakers were also not annotated. Furthermore, in contrast to Abe’s dataset, when the speaker was not on the page, such as texts from phone conversations or letters, we did not annotate them even if we knew the speaker. More details and examples of annotation rules are shown in the supplementary material.
(a) Abe’s annotations (character-to-text).
(b) Manga109Dialog (speaker-to-text).
Figure 2: Our dataset allows us to accurately determine whether the text on the right side is spoken by the character in the first frame or the character in the second frame, which Abe’s annotations were unable to do. ©Arai Satoshi.
Table 1: Statistics for two datasets.
Speaker-to-Text pairs
Dataset Annotated images Texts Easy Hard Total Pairs / page
Manga109Dialog 9,904 147,887 111,959 20,733 132,692 6.70
Abe’s annotations 9,830 147,887 - - 121,291 6.17
3.2 Data Analysis
(a) Easy.
(b) Hard.
(c) Total.
Figure 3: An example of annotations on three prediction difficulties. ©Kurita Riku.
We outsourced the annotation of our dataset to a company with expert annotators. After they completed the task, the first author confirmed and corrected their annotations. Creating the annotations took approximately three months. An overview of Manga109Dialog and Abe’s annotation dataset is shown in Table 1. In order to compare the two datasets at the same level, we did an additional step to determine the bounding boxes of speakers. For cases where the same character appears multiple times on a page, we preferred the box in the same frame as the text; otherwise, we would choose the nearest one. From Table 1, we can find Abe annotated 9,830 images of Manga109, containing an average of 6.17 speaker-to-text pairs per page, whereas we annotated all pages. In Manga109Dialog, 9,904 images include speaker-to-text pairs, with 6.70 pairs per page.
Moreover, we divided our annotations into Easy and Hard by the difficulty of prediction. If the speaker is in the same frame as the text, the text is considered Easy; otherwise, it is considered Hard. Total contains all pairs of speakers and texts. An example of annotations of varying difficulties is given in Figure 3. We performed experiments under these three difficulties to better evaluate our dataset in Section 5.3.
4 Approach
Here, we present our proposed scene-graph-based approach for comics speaker prediction using Manga109Dialog.
4.1 Problem setting
Let us first define our problem setting. Given an input image $I$, our task is twofold. First, we localize character regions and text regions. Next, for all combinations of characters and texts, we calculate the relationship score of each combination. The character with a higher score is more likely to be the speaker of the text. An example is shown in Figure 4. We represent each region as a bounding box $\mathbf{b}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$. For each region, we predict an object label $l\in\{\mathtt{character},\mathtt{text},\mathtt{background}\}$. For the combination of character $i$ and text $j$, the relationship score between $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ is represented as $r_{i\to j}\in\mathbb{R}$. This problem setting is equivalent to a standard SGG task, where the number of object classes is two (character or text) and the number of relationship labels is one (speak). In addition, we provide a frame detector and a frame order estimator. We can predict a bounding box of a frame; $\mathbf{p}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$. We can also obtain frame reading orders. We later use this information to enhance our system.
4.2 Speaker Detection via SGG
Introducing our approach in detail, we follow the pipeline of MOTIFS zellers2018neural , the most representative SGG model. Its process can be divided into three stages. The first stage is object detection. Given an image $I$, an object detector $f_{\mathrm{detect}}$ outputs a set of $N$ tuples.
$\{(\mathbf{b}_{i},\mathbf{f}_{i},\mathbf{l}_{i})\}_{i=1}^{N}=f_{\mathrm{detect}}(I).$ (1)
Each tuple consists of a bounding box $\mathbf{b}_{i}\in\mathbb{R}^{4}$, a feature vector $\mathbf{f}_{i}\in\mathbb{R}^{4096}$, and a vector of class label probabilities $\mathbf{l}_{i}\in\mathbb{R}^{3}$. The feature vector $\mathbf{f}_{i}$ encodes the visual information of the region $\mathbf{b}_{i}$. $\mathbf{l}_{i}$ represents the probability that the label is $\{\mathtt{character},\mathtt{text},\mathtt{background}\}$ respectively. Unlike MOTIFS, we use a Faster-RCNN ren2015faster with a ResNeXt-101-FPN lin2017feature ; xie2017aggregated backbone for this step.
Next, we fuse the detected feature vectors to produce a richer representation. We feed the set of the feature vectors and label probability vectors $\{(\mathbf{f}_{i},\mathbf{l}_{i})\}_{i=1}^{N}$ into the fusion module $f_{\mathrm{fuse}}$ to output enhanced features $\{\mathbf{d}_{i}\}_{i=1}^{N}$.
$\{\mathbf{d}_{i}\}_{i=1}^{N}=f_{\mathrm{fuse}}(\{(\mathbf{f}_{i},\mathbf{l}_{i})\}_{i=1}^{N}).$ (2)
Here, $\mathbf{d}_{i}\in\mathbb{R}^{512}$ is an enhanced representation for $\mathbf{b}_{i}$. Following MOTIFS, we adopt bidirectional LSTM models hochreiter1997long for this step. An additional LSTM is used to predict the object label $l_{i}$ from $\mathbf{l}_{i}$.
In the last stage, we predict the relationship score for each combination of all $N$ bounding boxes. Therefore, there are $N^{2}$ types of predictions. For example, the score of $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ ($r_{i\to j}\in\mathbb{R}$) is computed as
$r_{i\to j}=w(i,j)g(\mathbf{d}_{i},\mathbf{d}_{j},\mathbf{f}_{i,j}).$ (3)
$\mathbf{f}_{i,j}=f_{\mathrm{extract}}(\mathbf{b}_{i},\mathbf{b}_{j},\mathbf{f}_{i},\mathbf{f}_{j}).$ (4)
The function $g$ inputs the RoI feature of the union box $\textbf{f}_{i,j}$ and enhanced features $\mathbf{d}_{i},\mathbf{d}_{j}$ obtained from the previous stage and outputs a score. Here, $g$ consists of simple learnable matrices and a Softmax predicate classifier. Note that the feature vector $\textbf{f}_{i,j}$ is extracted from bounding boxes $\mathbf{b}_{i},\mathbf{b}_{j}$ and their feature vectors $\mathbf{f}_{i},\mathbf{f}_{j}$. The weight function $w(i,j)\in\mathbb{R}$ can be any function. If $w(i,j)=1$, the entire pipeline is the same as that of MOTIFS.
Figure 4: Framework of the proposed method. ©Arai Satoshi.
4.3 Use of Reading Order
Due to the unique feature of comics that the speaker is more likely to appear in the same or next frame that the text belongs to, we add frame information to the SGG model. To do so, we train another Faster-RCNN model to detect the frame regions from the input image. Following Kovanen and Ikuta’s study kovanen2015layered ; ikuta2023 , we can then estimate the reading order of the frames. Given a set of frame bounding boxes, we first decide whether they are horizontally divisible into two parts. If so, we split them and repeat the first step for each set until we cannot horizontally split them. Then, we decide whether each set is vertically divisible into two parts. If so, we do so and repeat the first step. When the set of frames cannot be divided horizontally and vertically, we number the frames according to reading order (from top to bottom, from right to left). As shown in Figure 4, we can obtain the result as the form of $\mathbf{p}_{1},\mathbf{p}_{2},\dots$, where $\mathbf{p}_{i}\in\mathbb{R}^{4}$ is a bounding box of the $i$-th frame. We also show an another example of frame detection and reading order estimation results in the supplementary material.
For each object, we calculate to which frame it belongs and obtain the reading order. We empirically found that the model performed best when the weight function $w(i,j)$ was in the form of Eq. 5:
$w(i,j)=\frac{1}{2+|k_{i}-k_{j}|}.$ (5)
Here, $k_{i}$ represents the reading order of the frame $\mathbf{p}$ to which $\mathbf{b}_{i}$ belongs. The closer the reading orders of $\mathbf{b}_{i}$, $\mathbf{b}_{j}$ are, the higher the score of $w(i,j)$. Therefore, we can make effect the prediction output by introducing frame orders and changing relationship scores.
To summarize, the end-to-end pipeline we designed is as follows. Given $I$, we obtain $N$ tuples $(\mathbf{b}_{i},\mathbf{f}_{i},\mathbf{l}_{i},\mathbf{d}_{i})$ by running Eq. 1 and 2. In addition, we obtain frame information $k_{i}$ through our frame detector. We then compute $r_{i\to j}$ for all $N^{2}$ possible combinations by Eq. 3. From the $N^{2}$ scores, we select the top $K$ items. Here, $K$ is a hyper-parameter. We discuss how to select $K$ in Section 4.4. After this, the model can generate a scene graph, consisting of object labels, their bounding boxes, and speaker-to-text pairs.
4.4 Evaluation Metrics
The earliest and the most widely accepted evaluation metric of SGG models is Recall@K lu2016visual . To calculate this, we first need to compose the set of subject-predicate-object into a triplet. For each triplet, we compute the prediction score by multiplying each score. We sort the scores and select the top $K$ triplets. We then compute the triplet recalls. We consider the triplet a correct prediction when all three elements are labeled correctly.
In classic SGG tasks, $K$ is usually predefined as 20, 50, or 100. However, since the number of texts fluctuates significantly from image to image, using a fixed $K$ value does not provide a fair comparison. Take Figure 5 as an example. When a large $K$ is fixed, Recall@K can cover almost all major combinations of objects if the number of texts is small, which makes the accuracy very close to 100% (left side of Figure 5). When a small $K$ is fixed, the score is always low if there are many texts (right side in Figure 5).
To compare our proposed method with rule-based methods more appropriately, we set $K$= #text and introduce a new evaluation metric called Recall@(#text), which is the Recall of selected predictions that can cover all texts on the page. For each text, we choose the triplet containing it with the highest score as the prediction. For the cases where there exist $N$ speakers for a single text, we select the top $N$ triplets as predictions. We use Recall@(#text) to evaluate the experimental results in Section 5.3.
Figure 5: Comparison of using Recall@K and Recall@(#text). ©Arai Satoshi. ©Akamatsu Ken.
5 Experiments
5.1 Experimental Setup
Baseline:
The simplest rule-based method is “shortest distance”, where we assume the text is spoken by the closest character. A slightly modified version is described as “frame distance”. We prioritize characters in the same frame as the text. If there are no characters in the frame, we make predictions using the same rule as “shortest distance”. In these two approaches, the amount of predictions is the same as that of texts. Therefore, there is no need to select the top $K$ predictions. Besides, we proposed a deep learning-based baseline, where the relationship score was calculated only by feeding the union feature $\textbf{f}_{i,j}$ obtained from Eq. 4 into a fully connected layer.
Datasets:
We divided Manga109Dialog into training and test sets with ratios of 70% and 30%, respectively. Furthermore, we divided the annotations into three levels of difficulty mentioned in Section 3.2. Besides, we tested our model on Abe’s dataset. Since the annotation in Abe’s dataset could not identify the particular bounding box of the speaker, we did an additional step and determined it by “frame distance”.
Evaluation:
In the task of SGG, the performance of the models can be evaluated in three protocols. (1) Predicate Classification (PredCls): predicting the relationship between two objects ($\{r_{i\to j}\}$), given an image ($I)$, object bounding boxes ($\{\mathbf{b}_{i}\}$), and object labels ($\{l_{i}\}$). (2) Scene Graph Classification (SGCls): predicting object labels and relationships, given an image and bounding boxes. (3) Scene Graph Detection (SGDet): detecting object bounding boxes and predicting their labels and relationships to generate an entire scene graph using only the given image. We executed the model for these three tasks to evaluate our proposed method.
5.2 Implementation Details
Because object detection is the most time-consuming step in speaker prediction, we pre-trained a Faster-RCNN model as our object detector. Following the previous studies in SGG tang2020unbiased ; suhail2021energy , we applied the Faster-RCNN with a ResNeXt-101-FPN backbone. We then froze the model as the object detector of our SGG models. Additionally, we pre-trained another Faster-RCNN as the frame detector. The Mean Average Precision (mAP) of the two models reached 86.09% and 96.48%, respectively. We present the details and results of our pre-trained object detector and frame detector in our supplementary material.
We trained our models on a single NVIDIA A100 GPU (60 GiB), with a batch size of 4. We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$. In the stage of SGG, we trained our SGG model on a single NVIDIA A100 GPU (60 GiB), with a batch size of 4. We optimized the model using SGD with an initial learning rate of $4\times 10^{-2}$. Our loss was the sum of the cross entropy for objects and relationships.
Figure 6: Recall@(#text) for PredCls.
Manga109Dialog
Method Easy Hard Total
Rule-based
    Short distance 71.43 22.72 63.41
    Frame distance 81.55 22.06 71.53
Deep learning-based
    Only w/ union feature 83.20 28.50 73.99
    Proposed w/o frame 84.73 29.55 75.48
    Proposed w/ frame 84.77 30.65 75.69
5.3 Quantitative Results
We ran our model under three tasks (PredCls, SGCls, SGDet). Because our study focuses on speaker prediction, we only show the results under PredCls in Table 6, meaning that the model only needs to find speaker-to-text pairs. See the supplementary material for the results under SGCls and SGDet. From Table 6, we can observe a significant improvement in our approach compared to the rule-based approach, especially on Hard. That was because the speaker in Hard is not in the same frame as the text, while “frame distance” gives preference to the character in the same frame. Therefore, other characters are in the frame to which the text belongs, “frame distance” is 100% likely to be wrong. Besides, the model using reading order information performed better than that without it.
In addition to our dataset, we also tested our model on Abe’s dataset under the task of PredCls. The results can be find in our supplementary material.
5.4 Qualitative Results
To better understand the generated scene graphs visually, we visualized the predictions. Figure 7 shows Recall@(#text) under PredCls. The green lines represent correct predictions, and the red lines represent incorrect predictions. On Easy, our model outperformed the “frame distance” method regardless of whether frame reading orders were used. For texts in Easy, if the speaker was not the closest character to the text, rule-based methods are certain to make an incorrect prediction, as shown in the red lines. Moreover, the results on Hard show that the proposed method of introducing frame orders exhibited a high accuracy in prediction and was able to handle more challenging cases. We show some additional examples in the supplementary material.
5.5 Challenges and Future Work
Since speaker prediction is somewhat subjective, more annotators are needed to enhance the reliability and validity of Manga109Dialog and enable it to be used more effectively for research and development in the field of comics analysis.
Besides, although our proposed method demonstrated significant improvements compared to conventional methods, there were still some cases where predictions failed. Two situations were primarily challenging for the machine. The first case was that the speaker was not in the closest position to the text. Though our model can cope with this situation to some degree, it was not always able to make correct predictions. The other case was that texts spoken by different characters appear alternately. Because no visual information can be used for prediction, this may be relatively difficult even for humans. The straightforward way to handle these cases would be to introduce Natural Language Processing (NLP) models. However, this research aimed to show to what extent speaker detection can be achieved only with visual information. The results of the present study provide a baseline needed to incorporate NLP in future research.
6 Conclusion
(a) Predictions on Easy. ©Gasan, ©Saijo Shinji.
(b) Predictions on Hard. ©Saki Kaori.
Figure 7: Examples of predictions made by rule-based method and proposed method. The green lines represent correct predictions, while the red lines represent incorrect predictions.
In this study, we have constructed a large-scale dialogue dataset called Manga109Dialog based on the shortcomings of existing annotations. we have presented a novel approach that applies SGG models to comics speaker detection, thereby creating a benchmark for deep learning-based methods where none existed before. Due to comics’ unique features, we have introduced frame reading order to help predict the speaker. To properly evaluate the experimental results, we have further proposed a new evaluation metric and tested our model on different test sets. The results indicate the reliability of Manga109Dialog, as well as the superior performance of our deep learning-based approach over existing rule-based methods. These findings highlight the potential of scene-graph-based methods in this field of comics speaker detection. We believe our dataset and model will have broader applicability in the digital processing of comic. Our work lays the groundwork for further investigations into the effective combination of SGG models and other techniques such as NLP, offering new insights for future research in this area.

### v2

@@
 Abstract
-The expanding market for e-comics has spurred interest in the development of automated methods to analyze comics.
-For further understanding of comics, an automated approach is needed to link text in comics to characters speaking the words.
-Comics speaker detection research has practical applications, such as automatic character assignment for audiobooks, automatic translation according to characters’ personalities, and inference of character relationships and stories.
-To deal with the problem of insufficient speaker-to-text annotations, we created a new annotation dataset Manga109Dialog [^1] Dataset and code are available at https://github.com/manga109/public-annotations.
-based on Manga109.
-Manga109Dialog is the world’s largest comics speaker annotation dataset, containing 132,692 speaker-to-text pairs.
-We further divided our dataset into different levels by prediction difficulties to evaluate speaker detection methods more appropriately.
-Unlike existing methods mainly based on distances, we propose a deep learning-based method using scene graph generation models.
-Due to the unique features of comics, we enhance the performance of our proposed model by considering the frame reading order.
-We conducted experiments using Manga109Dialog and other datasets.
-Experimental results demonstrate that our scene-graph-based approach outperforms existing methods, achieving a prediction accuracy of over 75%.
+The expanding market for e-comics has driven the development of automated methods for analyzing comics.
+To enhance the machine’s understanding of comics, an automated method is essential for linking text in comics to characters that speak those words.
+In this study, we developed Manga109Dialog [^1] Dataset and code are available at https://github.com/liyingxuan1012/Manga109Dialog., which is the world’s largest speaker-to-text annotation dataset for comics, containing 132,692 pairs.
+We proposed a novel deep learning-based method using scene graph generation models.
+To tailor the unique features of comics, we enhanced the performance by considering the frame reading order.
+Our experiments with Manga109Dialog show that our scene-graph-based approach outperforms existing methods, achieving a prediction accuracy of over 75%, thus establishing a robust benchmark for speaker detection in comics.
+Index Terms—  Dataset, Comic, Scene graph generation
 1 Introduction
 The market for e-comics has been expanding rapidly with the popularization of digital devices.
-In 2022, e-comics made up 66.2% of the Japanese comics market shuppankagaku2023 .
+In 2022, e-comics made up 66.2% of the Japanese comics market [1].
 The development of e-comics has led to increased interest in automated methods to analyses comics.
@@
 Additionally, automated methods to detect characters to whom text is attributed are necessary for effective speaker detection.
-Figure 1: An example of speaker detection. ©Akamatsu Ken.
-An example of speaker detection is shown in Figure 1.
+Fig. 1: An example of speaker detection. Courtesy of Akamatsu Ken.
+An example of speaker detection is shown in Fig.
+1.
 Given an image from comics, the system can detect characters, text, and frame regions, and can then automatically predict the character to which the text is attributed.
-This research can be applied to various tasks, such as automatic character assignment for text-to-speech reading, automatic translation according to characters’ personalities, inference of character relationships and stories, and automated generation of scenarios.
-Due to the lack of dialogue datasets and some limitations of existing annotations, we constructed more accurate and rich annotations, Manga109Dialog.
+This research can be applied to various tasks, such as automatic character assignment for text-to-speech reading, automatic translation according to characters’ personalities, inference of character relationships and stories.
+To address the significant gap in reliable speaker-to-text annotation datasets for comics, we developed Manga109Dialog.
 To the best of our knowledge, this is the largest comics dialogue dataset ever created.
-We divided the test set into two subsets based on the prediction difficulty.
-To evaluate the reliability and accuracy of Manga109Dialog, we conducted speaker detection experiments on it.
-Previous approaches were primarily based on the rule that the character nearest to a given text is most likely to be the speaker rigaud2015speech ; abe2020 .
-However, due to the unique and complex structure of comics, they may have difficulty making correct predictions in some complex cases.
-As an example, consider the text at bottom right in Figure 1.
+In order to validate the reliability and accuracy of Manga109Dialog, we conducted comprehensive speaker detection experiments on it.
+Since previous approaches were primarily based on the rule that the character nearest to a given text is most likely to be the speaker [2, 3], they may have difficulty making correct predictions in some complex cases.
+As an example, consider the text at bottom right in Fig.
+1.
 If we chose the character nearest to the text as the speaker, the boy would be incorrectly predicted as the speaker rather than the girl.
 Therefore, the relationship between characters and text should be considered to make a correct prediction.
-To address this challenge, we propose an innovative deep learning-based approach.
-We leverage scene graphs, which are widely recognized as one of the most popular methods for describing visual relationships.
-Scene graph generation (SGG) is the task of detecting objects and their relationships in an image.
-Since SGG models have performed well on real-world datasets, such as the Visual Genome dataset krishna2017visual , we propose utilizing SGG models for the task of comics speaker detection.
-Along with the general framework of SGG models, we introduce frame information because we find that the speaker tends to appear in the same frame as the text.
+To address this challenge, we propose a novel deep learning-based approach.
+We leverage scene graphs, a powerful tool for describing visual relationships.
+The task of scene graph generation (SGG) involves detecting objects and their relationships within an image.
+Given the successful application of SGG models in analyzing real-world datasets, such as the Visual Genome dataset [4], we propose their adaptation for comics speaker detection.
+In addition to employing the standard framework of SGG models, we incorporate frame information, as we find that the speaker tends to appear in the same frame as the text.
 Traditional metrics used to evaluate SGG lack suitability because the amount of speaker-to-text pairs varies from page to page.
-Besides, prediction results using traditional metrics cannot be compared to those of rule-based methods.
-Therefore, we present an evaluation metric well suited for this specific task, and demonstrate its superiority over conventional approaches in experimental results on Manga109Dialog.
+Therefore, we present an evaluation metric well suited for this specific task.
 The contributions are summarized as follows.
-  • We constructed an annotation dataset of associations between speakers and texts. This is the largest comics dialogue dataset in the world, containing over 130,000 speaker-to-text pairs.
-  • We propose a deep learning-based approach for comics speaker detection using SGG models. We enhance the results by introducing frame information in the relationship prediction stage.
-  • We introduce a new evaluation metric and a new standard for data division, that is, dividing the dataset according to prediction difficulty.
-Experimental results showed that Manga109Dialog provides a challenging yet realistic benchmark for comics speaker detection, and that our proposed approach exhibited 5% better performance compared to conventional rule-based methods.
-This research aimed to show to what extent speaker detection can be achieved solely with visual information, providing a benchmark for incorporating NLP in future research.
+  • We constructed Manga109Dialog, an annotation dataset of associations between speakers and texts. This is the largest comics dialogue dataset in the world.
+  • We proposed a deep learning-based approach for comics speaker detection using SGG models. We enhance the results by introducing frame information in the relationship prediction stage.
+  • We established a new benchmark for evaluation, setting a standard for future research in this domain.
+Experimental results demonstrated that Manga109Dialog provides a challenging yet realistic benchmark for comics speaker detection.
+Moreover, our proposed approach outperformed conventional rule-based methods by 5%, highlighting its effectiveness in complex cases.
 2 Related Work
 2.1 Comics Speaker Dataset
-Among the many comics available in different forms, only a few can be used for academic research due to copyright issues.
-eBDtheque guerin2013ebdtheque is a dataset of comics available for direct use.
-It comprises 100 images from French, American, and Japanese comics, containing 1,550 characters, 1,092 speech balloons, and 887 speaker-to-text pairs.
-Moreover, a large-scale manga dataset called Manga109 aizawa2020building contains 109 Japanese comics, including 21,142 images, 2,979 characters, and 147,918 texts.
-Given the lack of available comics datasets associated with speakers and comics texts, Abe et al.
-abe2020 constructed a dataset based on Manga109 with a total of 121,364 annotated speaker-to-text pairs.
+Given the scarcity of available comics speaker datasets, Abe et al.
+[3] first constructed a dataset based on Manga109 [5], which contains 109 Japanese comics and annotations for both character and text regions.
+Although Abe’s work is pioneering, the dataset has some notable limitations.
+First, it does not have explicit links between the text and the character in the comics – it only adds the character name to the text without specifying the bounding box of the character, which poses a challenge when the same character appears multiple times on a page.
+Additionally, the tendency to annotate simpler relationships results in exclusion of more complex cases.
+To address these shortcomings, we developed Manga109Dialog, employing a unique annotation standard to provide a more comprehensive and applicable dataset.
 2.2 Comics Speaker Detection
-Rule-Based Methods:
-Rigaud et al.
-rigaud2015speech proposed early rule-based speaker prediction method using distance calculation between speech balloon centroids and character centroids.
-The character with the shortest distance was assumed to be the speaker.
-They performed experiments on eBDtheque and obtained an accuracy of 78.58%.
-In the study by Abe et al.
-on Manga109, four types of information were used abe2020 , including the text-character distance, the characters in the same frame with the text, the direction of the tail of the speech balloon, and language style (such as first-person utterances).
-Their approach experimentally obtained an accuracy of 70.7%.
-Deep Learning-Based Methods:
-In recent years, deep learning-based methods have also been developed.
-Yamamoto et al.
-yamamoto2018 used features extracted from the input images and metadata to calculate scores for each text-character pair, while other studies explored time-series learning nakagawa2019 and Natural Language Processing yamaguchi2021 .
-The results of these studies on Manga109 all demonstrate the potential of learning-based methods in speaker prediction.
-However, their results were only based on a part of the dataset and lacked implementation details.
-Therefore, a large dataset is needed to evaluate these models under a unified standard.
+Traditional approaches in comics speaker prediction typically focus on predicting the relationships between characters and text without detecting specific object regions.
+These methods, primarily rule-based [2, 3], determine the speaker by calculating the distances between text and characters, which often makes them struggle in complex cases.
+Recently, there has been a shift towards deep learning-based approaches, such as extracting features from objects or employing time-series learning techniques [6, 7].
+Although they demonstrated the potential of deep learning-based approaches, they were limited by using only partial datasets and showed suboptimal performance.
+This gap underscores the need for a complete dataset and a unified standard for evaluation.
+In our study, we introduce object detection into the speaker prediction evaluation, aligning more closely with real-world applications.
+We also develop a comprehensive benchmark for this task.
 2.3 Scene Graph Generation (SGG)
-Scene graphs were introduced by Johnson et al.
-johnson2015image as a data structure for describing objects in a scene.
-To generate a scene graph, an SGG model first detects object regions and their categories, then identifies relationships between objects.
+Scene graph is a data structure for describing objects in a scene.
 Since scene graphs provide rich semantic features, they can distinguish images or videos more accurately and describe them more precisely.
-Scene graphs have demonstrated success in various visual tasks, including image retrieval, image captioning, and visual question answering johnson2015image ; yao2018exploring ; tang2019learning .
+Scene graphs have demonstrated success in various visual tasks, including image retrieval, image captioning, and visual question answering [8, 9, 10].
 3 Manga109Dialog
-Abe et al.
-abe2020 first created a comics speaker dataset based on the entire Manga109 dataset.
-Although their study was pioneering and impressive, this dataset still involves some notable limitations.
-First, they linked the name of the character to the text instead of the bounding box of the speaker.
-Therefore, we cannot specify a specific bounding box if the same character appears multiple times on a single page.
-Besides, since annotators prefer making annotations of simple relationships, some more complicated cases are not included in this dataset.
-Aiming to produce a more complete and applicable dataset for speaker detection in comics, we constructed a new dataset, Manga109Dialog.
+Fig. 2: Annotation rules. Courtesy of Inokuma Shinobu.
+3.1 Dataset Construction
+In this section, we show the details of how we constructed Manga109Dialog.
+Aiming to produce a more complete and applicable dataset for speaker detection in comics, we created our dataset in a distinct annotation standard.
 Instead of connecting the character name to the text, we connect the bounding box of the speaker to the text.
-The visual differences between Abe’s dataset and Manga109Dialog are shown as Figure 2.
-3.1 Dataset Construction
-In this section, we show the details of how we constructed our dataset. Our annotations were mainly based on the following rules.
+The annotations were mainly based on the following rules.
   • If the character to whom the text is attributed appears only one on the page, link the bounding box of the character to the text regardless of the position.
   • If more than one character is on the page, the bounding box of the speaker in the same frame as the text is linked in priority.
-If the speaker is not in the same frame, it is determined by the reading order (from top to bottom, right to left).
-The texts in the last frame of the page, however, are linked to the character in the second-to-last frame.
+If the speaker is not in the same frame, it is determined by the reading order (from top to bottom, right to left), as shown in Fig.
+2.
   • When there is more than one speaker, we link the text to all speakers.
   • Texts spoken by “Others” are annotated if we can specify the bounding box of the speaker, and are excluded otherwise.
-Although we make annotations on Manga109, not all texts are our annotation targets.
-For example, titles and descriptions were not annotated.
-Texts designed as “Narration” or with unknown speakers were also not annotated.
-Furthermore, in contrast to Abe’s dataset, when the speaker was not on the page, such as texts from phone conversations or letters, we did not annotate them even if we knew the speaker.
 More details and examples of annotation rules are shown in the supplementary material.
-(a) Abe’s annotations (character-to-text).
-(b) Manga109Dialog (speaker-to-text).
-Figure 2: Our dataset allows us to accurately determine whether the text on the right side is spoken by the character in the first frame or the character in the second frame, which Abe’s annotations were unable to do.
-©Arai Satoshi.
-Table 1: Statistics for two datasets.
+Table 1: An overview of Manga109Dialog.
+Item Number of elements
+Images 10,602
+Annotated Images 9,904
+Texts 147,887
 Speaker-to-Text Pairs
-Dataset Annotated images Texts Easy Hard Total Pairs / page
-Manga109Dialog 9,904 147,887 111,959 20,733 132,692 6.70
-Abe’s annotations 9,830 147,887 - - 121,291 6.17
+      Easy 111,959
+      Hard 20,733
+      Total 132,692
+Pairs / Page (Total) 6.70
 3.2 Data Analysis
-(a) Easy.
-(b) Hard.
-(c) Total.
-Figure 3: An example of annotations on three prediction difficulties. ©Kurita Riku.
-We outsourced the annotation of our dataset to a company with expert annotators.
-After they completed the task, the first author confirmed and corrected their annotations.
-Creating the annotations took approximately three months.
-An overview of Manga109Dialog and Abe’s annotation dataset is shown in Table 1.
-In order to compare the two datasets at the same level, we did an additional step to determine the bounding boxes of speakers.
-For cases where the same character appears multiple times on a page, we preferred the box in the same frame as the text; otherwise, we would choose the nearest one.
-From Table 1, we can find Abe annotated 9,830 images of Manga109, containing an average of 6.17 speaker-to-text pairs per page, whereas we annotated all pages.
-In Manga109Dialog, 9,904 images include speaker-to-text pairs, with 6.70 pairs per page.
-Moreover, we divided our annotations into Easy and Hard by the difficulty of prediction.
+Creating the annotations for Manga109Dialog took approximately three months.
+An overview of the dataset is presented in Table 1.
+Manga109Dialog comprises 9,904 images featuring speaker-to-text pairs, averaging 6.70 pairs per page.
+We further divided our annotations into Easy and Hard by prediction difficulties.
 If the speaker is in the same frame as the text, the text is considered Easy; otherwise, it is considered Hard.
 Total contains all pairs of speakers and texts.
-An example of annotations of varying difficulties is given in Figure 3.
-We performed experiments under these three difficulties to better evaluate our dataset in Section 5.3.
+An example of annotations of varying difficulties is given in Fig.
+3.
+We conducted experiments under these three difficulties to evaluate speaker detection methods more appropriately.
 4 Approach
-Here, we present our proposed scene-graph-based approach for comics speaker prediction using Manga109Dialog.
 4.1 Problem Setting
-Let us first define our problem setting.
+Before introducing our proposed scene-graph-based approach, let us first define our problem setting.
+Unlike existing studies on comics speaker prediction, we incorporate the steps for detecting and identifying object regions.
 Given an input image $I$, our task is twofold.
 First, we localize character regions and text regions.
-Next, for all combinations of characters and texts, we calculate the relationship score of each combination.
-The character with a higher score is more likely to be the speaker of the text.
-An example is shown in Figure 4.
+Then, for all combinations of characters and texts, we calculate the relationship score of each combination.
+The character with a higher score is more likely to be the speaker.
+An example of the pipeline is shown in Fig.
+4.
 We represent each region as a bounding box $\mathbf{b}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$.
@@
 This problem setting is equivalent to a standard SGG task, where the number of object classes is two (character or text) and the number of relationship labels is one (speak).
-In addition, we provide a frame detector and a frame order estimator.
-We can predict a bounding box of a frame; $\mathbf{p}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$.
-We can also obtain frame reading orders.
+In addition, we develop a frame detector and a frame order estimator to detect frame bounding boxes ($\mathbf{p}=[x,y,w,h]^{\top}\in\mathbb{R}^{4}$) and determine their reading order.
 We later use this information to enhance our system.
+(a) Easy.
+(b) Hard.
+(c) Total.
+Fig. 3: An example of annotations on three prediction difficulties. Courtesy of Kurita Riku.
+Fig. 4: Framework of the proposed method. Courtesy of Arai Satoshi.
 4.2 Speaker Detection via SGG
-Introducing our approach in detail, we follow the pipeline of MOTIFS zellers2018neural , the most representative SGG model.
-Its process can be divided into three stages.
+Our approach is based on the MOTIFS pipeline [11, 12], the most representative SGG model.
+Its process is divided into three stages.
 The first stage is object detection.
@@
 Each tuple consists of a bounding box $\mathbf{b}_{i}\in\mathbb{R}^{4}$, a feature vector $\mathbf{f}_{i}\in\mathbb{R}^{4096}$, and a vector of class label probabilities $\mathbf{l}_{i}\in\mathbb{R}^{3}$.
-The feature vector $\mathbf{f}_{i}$ encodes the visual information of the region $\mathbf{b}_{i}$.
-$\mathbf{l}_{i}$ represents the probability that the label is $\{\mathtt{character},\mathtt{text},\mathtt{background}\}$ respectively.
-Unlike MOTIFS, we use a Faster-RCNN ren2015faster with a ResNeXt-101-FPN lin2017feature ; xie2017aggregated backbone for this step.
+Unlike MOTIFS, we employ a Faster-RCNN [13] with a ResNeXt-101-FPN [14, 15] backbone for this step.
 Next, we fuse the detected feature vectors to produce a richer representation.
@@
 Here, $\mathbf{d}_{i}\in\mathbb{R}^{512}$ is an enhanced representation for $\mathbf{b}_{i}$.
-Following MOTIFS, we adopt bidirectional LSTM models hochreiter1997long for this step.
+We adopt bidirectional LSTM models [16] for this step.
 An additional LSTM is used to predict the object label $l_{i}$ from $\mathbf{l}_{i}$.
-In the last stage, we predict the relationship score for each combination of all $N$ bounding boxes.
-Therefore, there are $N^{2}$ types of predictions.
+In the last stage, we calculate the relationship score for each pair among the $N$ bounding boxes, resulting in $N^{2}$ possible combinations.
 For example, the score of $\mathbf{b}_{i}$ and $\mathbf{b}_{j}$ ($r_{i\to j}\in\mathbb{R}$) is computed as
@@
 $\mathbf{f}_{i,j}=f_{\mathrm{extract}}(\mathbf{b}_{i},\mathbf{b}_{j},\mathbf{f}_{i},\mathbf{f}_{j}).$ (4)
-The function $g$ inputs the RoI feature of the union box $\textbf{f}_{i,j}$ and enhanced features $\mathbf{d}_{i},\mathbf{d}_{j}$ obtained from the previous stage and outputs a score.
+The function $g$ inputs the union box’s RoI feature $\textbf{f}_{i,j}$ and enhanced features $\mathbf{d}_{i},\mathbf{d}_{j}$ and outputs a score.
 Here, $g$ consists of simple learnable matrices and a Softmax predicate classifier.
-Note that the feature vector $\textbf{f}_{i,j}$ is extracted from bounding boxes $\mathbf{b}_{i},\mathbf{b}_{j}$ and their feature vectors $\mathbf{f}_{i},\mathbf{f}_{j}$.
+The feature vector $\textbf{f}_{i,j}$ is extracted from bounding boxes $\mathbf{b}_{i},\mathbf{b}_{j}$ and their feature vectors $\mathbf{f}_{i},\mathbf{f}_{j}$.
 The weight function $w(i,j)\in\mathbb{R}$ can be any function.
 If $w(i,j)=1$, the entire pipeline is the same as that of MOTIFS.
-Figure 4: Framework of the proposed method. ©Arai Satoshi.
 4.3 Use of Reading Order
 Due to the unique feature of comics that the speaker is more likely to appear in the same or next frame that the text belongs to, we add frame information to the SGG model.
-To do so, we train another Faster-RCNN model to detect the frame regions from the input image.
-Following Kovanen and Ikuta’s study kovanen2015layered ; ikuta2023 , we can then estimate the reading order of the frames.
-Given a set of frame bounding boxes, we first decide whether they are horizontally divisible into two parts.
-If so, we split them and repeat the first step for each set until we cannot horizontally split them.
-Then, we decide whether each set is vertically divisible into two parts.
-If so, we do so and repeat the first step.
-When the set of frames cannot be divided horizontally and vertically, we number the frames according to reading order (from top to bottom, from right to left).
-As shown in Figure 4, we can obtain the result as the form of $\mathbf{p}_{1},\mathbf{p}_{2},\dots$, where $\mathbf{p}_{i}\in\mathbb{R}^{4}$ is a bounding box of the $i$-th frame.
-We also show an another example of frame detection and reading order estimation results in the supplementary material.
+We train another Faster-RCNN model to detect the frame regions from the input image.
+To estimate the reading order of these frames, we follow the algorithm described in Kovanen and Ikuta’s study [17, 18].
+As shown in Fig.
+4, the outcome is a series of ordered frame bounding boxes $\mathbf{p}_{1},\mathbf{p}_{2},\dots$, where $\mathbf{p}_{i}\in\mathbb{R}^{4}$ is a bounding box of the $i$-th frame.
+We show the details of this process and results of frame detection and reading order estimation in the supplementary material.
 For each object, we calculate to which frame it belongs and obtain the reading order. We empirically found that the model performed best when the weight function $w(i,j)$ was in the form of Eq. 5:
@@
 The closer the reading orders of $\mathbf{b}_{i}$, $\mathbf{b}_{j}$ are, the higher the score of $w(i,j)$.
-Therefore, we can make effect the prediction output by introducing frame orders and changing relationship scores.
-To summarize, the end-to-end pipeline we designed is as follows.
-Given $I$, we obtain $N$ tuples $(\mathbf{b}_{i},\mathbf{f}_{i},\mathbf{l}_{i},\mathbf{d}_{i})$ by running Eq.
-1 and 2.
-In addition, we obtain frame information $k_{i}$ through our frame detector.
-We then compute $r_{i\to j}$ for all $N^{2}$ possible combinations by Eq.
-3.
-From the $N^{2}$ scores, we select the top $K$ items.
-Here, $K$ is a hyper-parameter.
-We discuss how to select $K$ in Section 4.4.
-After this, the model can generate a scene graph, consisting of object labels, their bounding boxes, and speaker-to-text pairs.
+Therefore, we can enhance prediction accuracy by incorporating frame orders and adjusting relationship scores accordingly.
 4.4 Evaluation Metrics
-The earliest and the most widely accepted evaluation metric of SGG models is Recall@K lu2016visual .
-To calculate this, we first need to compose the set of subject-predicate-object into a triplet.
-For each triplet, we compute the prediction score by multiplying each score.
-We sort the scores and select the top $K$ triplets.
-We then compute the triplet recalls.
-We consider the triplet a correct prediction when all three elements are labeled correctly.
-In classic SGG tasks, $K$ is usually predefined as 20, 50, or 100.
-However, since the number of texts fluctuates significantly from image to image, using a fixed $K$ value does not provide a fair comparison.
-Take Figure 5 as an example.
-When a large $K$ is fixed, Recall@K can cover almost all major combinations of objects if the number of texts is small, which makes the accuracy very close to 100% (left side of Figure 5).
-When a small $K$ is fixed, the score is always low if there are many texts (right side in Figure 5).
-To compare our proposed method with rule-based methods more appropriately, we set $K$= #text and introduce a new evaluation metric called Recall@(#text), which is the Recall of selected predictions that can cover all texts on the page.
-For each text, we choose the triplet containing it with the highest score as the prediction.
-For the cases where there exist $N$ speakers for a single text, we select the top $N$ triplets as predictions.
-We use Recall@(#text) to evaluate the experimental results in Section 5.3.
-Figure 5: Comparison of using Recall@K and Recall@(#text). ©Arai Satoshi. ©Akamatsu Ken.
+The earliest and the most widely accepted evaluation metric for SGG models is Recall@K, traditionally predefined at 20, 50, or 100 [19].
+However, since the number of texts in comics fluctuates significantly from image to image, using a fixed $K$ value does not provide a fair comparison.
+Take Fig.
+5 as an example.
+When a large $K$ is fixed, Recall@K can cover almost all major combinations of objects if the number of texts is small, which makes the accuracy very close to 100% (left side of Fig.
+5).
+When a small $K$ is fixed, the score is always low if there are many texts (right side in Fig.
+5).
+To establish a more reliable benchmark for the task of comics speaker detection, we introduce a new evaluation metric called Recall@(#text), which measures the Recall of selected predictions covering all texts on a page.
+For each text, we choose the subject-predicate-object triplet containing it with the highest score as the prediction.
+In cases where a single text has $N$ speakers, we select the top $N$ triplets.
+We consider a triplet correct if the object regions are detected with an IoU greater than 0.65 and correctly labeled.
+Fig. 5: Comparison of using Recall@K and Recall@(#text). Courtesy of Arai Satoshi, Akamatsu Ken.
 5 Experiments
 5.1 Experimental Setup
-Baseline:
-The simplest rule-based method is “shortest distance”, where we assume the text is spoken by the closest character.
+Baseline: The simplest rule-based method is “shortest distance”, where we assume the text is spoken by the closest character.
 A slightly modified version is described as “frame distance”.
@@
 If there are no characters in the frame, we make predictions using the same rule as “shortest distance”.
-In these two approaches, the amount of predictions is the same as that of texts.
-Therefore, there is no need to select the top $K$ predictions.
-Besides, we proposed a deep learning-based baseline, where the relationship score was calculated only by feeding the union feature $\textbf{f}_{i,j}$ obtained from Eq.
+Besides, we propose a deep learning-based baseline, where the relationship score was calculated only by feeding the union feature $\textbf{f}_{i,j}$ obtained from Eq.
 4 into a fully connected layer.
-Datasets:
-We divided Manga109Dialog into training and test sets with ratios of 70% and 30%, respectively.
-Furthermore, we divided the annotations into three levels of difficulty mentioned in Section 3.2.
-Besides, we tested our model on Abe’s dataset.
-Since the annotation in Abe’s dataset could not identify the particular bounding box of the speaker, we did an additional step and determined it by “frame distance”.
-Evaluation:
-In the task of SGG, the performance of the models can be evaluated in three protocols.
-(1) Predicate Classification (PredCls): predicting the relationship between two objects ($\{r_{i\to j}\}$), given an image ($I)$, object bounding boxes ($\{\mathbf{b}_{i}\}$), and object labels ($\{l_{i}\}$).
+Datasets: We divided Manga109Dialog into training and test sets with ratios of 70% and 30%, respectively.
+Furthermore, to see our model’s behavior on different styles of comics or languages, we tested it on eBDtheque, which comprises 100 images from French, American, and Japanese comics [20].
+Evaluation: In the task of SGG, the performance of the models can be evaluated in three protocols.
+(1) Predicate Classification (PredCls): predicting the relationship between two objects, given an image $I$, object bounding boxes $\{\mathbf{b}_{i}\}$, and object labels $\{l_{i}\}$.
 (2) Scene Graph Classification (SGCls): predicting object labels and relationships, given an image and bounding boxes.
-(3) Scene Graph Detection (SGDet): detecting object bounding boxes and predicting their labels and relationships to generate an entire scene graph using only the given image.
+(3) Scene Graph Detection (SGDet): detecting object bounding boxes and predicting their labels and relationships using only the given image.
 We executed the model for these three tasks to evaluate our proposed method.
-5.2 Implementation Details
-Because object detection is the most time-consuming step in speaker prediction, we pre-trained a Faster-RCNN model as our object detector.
-Following the previous studies in SGG tang2020unbiased ; suhail2021energy , we applied the Faster-RCNN with a ResNeXt-101-FPN backbone.
-We then froze the model as the object detector of our SGG models.
+Implementation Details: Because object detection is the most time-consuming step, we pre-trained a Faster-RCNN model as our object detector.
 Additionally, we pre-trained another Faster-RCNN as the frame detector.
-The Mean Average Precision (mAP) of the two models reached 86.09% and 96.48%, respectively.
-We present the details and results of our pre-trained object detector and frame detector in our supplementary material.
-We trained our models on a single NVIDIA A100 GPU (60 GiB), with a batch size of 4.
-We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$.
-In the stage of SGG, we trained our SGG model on a single NVIDIA A100 GPU (60 GiB), with a batch size of 4.
-We optimized the model using SGD with an initial learning rate of $4\times 10^{-2}$.
-Our loss was the sum of the cross entropy for objects and relationships.
-Figure 6: Recall@(#text) for PredCls.
+The Mean Average Precision (mAP) of the two detectors reached 86.09% and 96.48%, respectively.
+We present the details of them in the supplementary material.
+In the stage of SGG, we trained our model on a single NVIDIA A100 GPU, using the SGD optimizaer with a batch size of 4.
+We used the WarmupReduceLROnPlateau as the learning scheduler and set the initial learning rate to $4\times 10^{-2}$.
+The loss was the sum of the cross entropy for objects and relationships.
+Table 2: Recall@(#text) for PredCls.
 Manga109Dialog
@@
 See the supplementary material for the results under SGCls and SGDet.
-From Table 6, we can observe a significant improvement in our approach compared to the rule-based approach, especially on Hard.
+Table 2 shows an improvement in our approach over rule-based approaches, especially on Hard.
 That was because the speaker in Hard is not in the same frame as the text, while “frame distance” gives preference to the character in the same frame.
@@
 Besides, the model using reading order information performed better than that without it.
-In addition to our dataset, we also tested our model on Abe’s dataset under the task of PredCls. The results can be find in our supplementary material.
+Additionally, we directly tested our model on eBDtheque.
+Our model achieved 77.45% accuracy under PredCls, slightly below “frame distance”.
+This is mainly due to style variations in old European and American comics, where the speaker and text are often in the same frame.
+More details are available in the supplementary material.
 5.3 Qualitative Results
-To better understand the generated scene graphs visually, we visualized the predictions.
-Figure 7 shows Recall@(#text) under PredCls.
+To better understand the generated scene graphs, we visualized the predictions.
+Fig.
+6 shows Recall@(#text) under PredCls.
 The green lines represent correct predictions, and the red lines represent incorrect predictions.
-On Easy, our model outperformed the “frame distance” method regardless of whether frame reading orders were used.
-For texts in Easy, if the speaker was not the closest character to the text, rule-based methods are certain to make an incorrect prediction, as shown in the red lines.
+For texts in Easy, if the speaker was not the closest character to the text, rule-based methods are certain to make an incorrect prediction, while our model outperformed the “frame distance” method regardless of whether frame reading orders were used.
 Moreover, the results on Hard show that the proposed method of introducing frame orders exhibited a high accuracy in prediction and was able to handle more challenging cases.
 We show some additional examples in the supplementary material.
+(a) Predictions on Easy.
+(b) Predictions on Hard.
+Fig.
+6: Examples of predictions made by rule-based method and proposed method.
+Courtesy of Gasan, Saijo Shinji, Saki Kaori.
+The green lines represent correct predictions, while the red lines represent incorrect predictions.
 5.4 Challenges and Future Work
-Since speaker prediction is somewhat subjective, more annotators are needed to enhance the reliability and validity of Manga109Dialog and enable it to be used more effectively for research and development in the field of comics analysis.
-Besides, although our proposed method demonstrated significant improvements compared to conventional methods, there were still some cases where predictions failed.
-Two situations were primarily challenging for the machine.
-The first case was that the speaker was not in the closest position to the text.
-Though our model can cope with this situation to some degree, it was not always able to make correct predictions.
-The other case was that texts spoken by different characters appear alternately.
+Although our proposed method demonstrated improvements compared to conventional methods, there were two situations primarily challenging.
+(1) The speaker was not in the closest position to the text.
+Though our model can cope with this situation to some degree, it did not always make correct predictions.
+(2) Texts spoken by different characters appear alternately.
 Because no visual information can be used for prediction, this may be relatively difficult even for humans.
@@
 6 Conclusion
-(a) Predictions on Easy. ©Gasan, ©Saijo Shinji.
-(b) Predictions on Hard. ©Saki Kaori.
-Figure 7: Examples of predictions made by rule-based method and proposed method. The green lines represent correct predictions, while the red lines represent incorrect predictions.
-In this study, we have constructed a large-scale dialogue dataset called Manga109Dialog based on the shortcomings of existing annotations.
-we have presented a novel approach that applies SGG models to comics speaker detection, thereby creating a benchmark for deep learning-based methods where none existed before.
-Due to comics’ unique features, we have introduced frame reading order to help predict the speaker.
-To properly evaluate the experimental results, we have further proposed a new evaluation metric and tested our model on different test sets.
-The results indicate the reliability of Manga109Dialog, as well as the superior performance of our deep learning-based approach over existing rule-based methods.
-These findings highlight the potential of scene-graph-based methods in this field of comics speaker detection.
-We believe our dataset and model will have broader applicability in the digital processing of comic.
-Our work lays the groundwork for further investigations into the effective combination of SGG models and other techniques such as NLP, offering new insights for future research in this area.
+In this study, we developed Manga109Dialog, a large-scale dialogue dataset that significantly advances the field of comics analysis.
+We pioneered a novel approach by applying SGG models to the task of comics speaker detection, establishing a benchmark for deep learning-based methods where none existed before.
+The results not only demonstrate the reliability of Manga109Dialog but also show the superior efficacy of our scene-graph-based approach.
+Our work paves the way for future research, particularly in exploring the effective combination of SGG models and other techniques such as NLP, offering new insights in this field.
+Appendix A Dataset Construction Details
+This section shows the construction details of Manga109Dialog.
+A.1 Annotation Rules
+In this section, we use examples to show more details of annotation rules stated in Section 3.1.
+First, let us introduce the basic rules.
+Although the images in Manga109 are double-page spreads, from the semantic point of view, we take every single page as a unit.
+If the number of the character who is saying the text is only one on the page, link the bounding box of the character to the text regardless of the position.
+If the number of the character who is saying the text is more than one on the page, the bounding box of the character in the same frame as the text is linked in priority.
+If the character is not in the same frame, it is determined by the reading order (from top to bottom, right to left).
+Take the rightmost text of Fig.
+A (a) as an example.
+Since there is no character in the frame with it, it is linked to the boy who appears in the middle frame according to the reading order.
+(a) If the speaker is not in the same frame, it is determined by the reading order.
+(b) If the texts are in the last frame of the page, they are linked to the character in the second-to-last frame.
+Fig. A: Examples where the speaker is not in the same frame as the text. Courtesy of Inokuma Shinobu, Kurita Riku.
+However, there exist exceptions.
+For texts in the last frame of the page, since we only consider the single page, we link the character in the second-to-last frame to the text instead of the character in the next page.
+One example is shown in Fig.
+A (b).
+Next, let us consider some special cases. Take Fig. B (a) as an example. When there is more than one speaker for a single text, we link the text to the bounding boxes of all speakers.
+Besides, for texts said by “Others”, if we can specify the bounding box of the speaker (such as Fig. B (b)), it is annotated; otherwise, it is excluded.
+Although we make annotations on Manga109, not all texts are our annotation targets.
+For example, titles and descriptions are not annotated.
+For texts belonging to “Narration” or with unknown speakers, like those in Fig.
+C (a), we do not annotate them, either.
+Furthermore, when the speaker is not on the page, such as texts from phone conversations (Fig.
+C (b)) or letters, the text is not in the scope of annotation.
+(a) An example of multiple speakers for a single text.
+(b) An example of texts said by “Others”.
+Fig. B: Some special cases. Courtesy of Akamatsu Ken, Kobayashi Yuki.
+(a) Texts with unknown speakers are not our annotation targets.
+(b) When the speaker is not on the page, the text is not annotated.
+Fig. C: Texts not in the scope of annotation. Courtesy of Inokuma Shinobu, Akamatsu Ken.
+A.2 Annotation Details
+We outsourced the annotation of Manga109Dialog to a company with expert annotators, consisting of 8 operators and 3 inspectors.
+Operators used the annotation tool shown in Fig.
+D to link the speaker to the text.
+Then inspectors checked the annotations for omissions and mistakes.
+Creating the annotations took approximately three months, and work hours (except for inspections) exceeded 627 hours.
+The annotations for each comic are saved in an xml file, an example is shown in Fig.
+E.
+A.3 Visualization Tool
+After the annotation company completed the task, we confirmed and corrected their annotations.
+To check the annotations in a visual way, we designed a visualization tool.
+The interface is shown in Fig.
+F.
+When pressing the “relationship” button, the visualization tool can show the regions of characters and texts with different colored rectangular boxes.
+Additionally, it can show characters’ names and speaker-to-text pairs.
+As shown in Fig.
+F, each orange line connects the center coordinates of the speaker and the text.
+When a text has $n$ speakers, there will be $n$ lines.
+Fig. D: The interface of the annotation tool. Courtesy of Yoshi Masako.
+Fig. E: Format of the saved annotations.
+Fig. F: The interface of our visualization tool. Courtesy of Kato Masaki.
+Appendix B Experimental Preparation
+This section visually explains the details of our pre-trained object detector, frame detector, and frame order estimator.
+(a) Ground truth.
+(b) Detection results.
+Fig. G: An example of detection results. Courtesy of Matsuda Naomasa.
+B.1 Object Detector
+We pre-trained our object detector to save run time of speaker prediction.
+Following the previous studies in SGG [12], we applied the Faster-RCNN with a ResNeXt-101-FPN backbone.
+We then froze the model as the object detector.
+We trained our models on a single NVIDIA A100 GPU, with a batch size of 4.
+We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$.
+We set object classes to three: character, text and background.
+The mAP of our detector was 86.09%.
+The high accuracy may be related to the small number of object categories in our experiments.
+An example of detection results is shown in Fig.
+G.
+The left side of Fig.
+G shows the visualization of ground truth, and the right side shows the detected bounding boxes and predicted labels.
+We can also see from this example that our model has a high accuracy.
+B.2 Frame Detector and Frame Order Estimator
+We pre-trained another Faster-RCNN model as the frame detector, which was trained on a single NVIDIA A100 GPU, with a batch size of 4.
+We used SGD as an optimizer and set the initial learning rate to $4\times 10^{-3}$.
+We set object classes to two: frame and background and our frame detector achieved an mAP of 96.48%.
+For detected frames, we used the algorithm described in Kovanen and Ikuta’s study [17, 18] to presume their reading orders.
+Given a set of frame bounding boxes, we first decide whether they are horizontally divisible into two parts.
+If so, we split them and repeat the first step for each set until we cannot horizontally split them.
+Then, we decide whether each set is vertically divisible into two parts.
+If so, we do so and repeat the first step.
+When the set of frames cannot be divided horizontally and vertically, we number the frames according to reading order (from top to bottom, from right to left).
+An example of the estimation results is shown in Fig.
+H.
+When there is a high degree of overlap in a set of frames, we consider the set inseparable.
+All frames in this set are given the same order and are shown in orange dotted-line bounding boxes.
+For the example in Fig.
+H, the first frame covers half of the page in ground truth.
+Since the following three frames are inside the first frame, the first four frames are considered inseparable and share the same reading order.
+Therefore, although the detected frame is different from those in ground truth, the reading order is presumed correctly, so there is no negative impact on our experiments of speaker prediction.
+(a) Ground truth.
+(b) Detection results.
+Fig. H: Results of frame detection and reading order estimation. Courtesy of Hikochi Sakuya.
+Appendix C Rule-Based Baseline
+Here, we recap simple rule-based approaches using our notations in Section 4.2.
+The simplest baseline is “shortest distance”. That is, the score $r_{i\to j}$ is defined as follows.
+$r_{i\to j}=\begin{cases}1&\mathrm{if}~{}\mathbf{b}_{i}~{}\mathrm{is~{}closest~{}to}~{}\mathbf{b}_{j},\\ &l_{i}=\mathtt{character},~{}~{}\mathrm{and}~{}~{}l_{j}=\mathtt{text}.\\ 0&\mathrm{else}.\end{cases}$ (a)
+Here, we assume that $\{\mathbf{b}_{i}\}_{i=1}^{N}$ is given by $f_{\mathrm{detect}}$.
+This equation means that the text is spoken by the closest character.
+This is simple and can be computed rapidly, but cannot handle complex cases.
+A slightly modified version is described as “frame distance”, which is defined as given below.
+[…]
+Here, we prioritize characters in the same frame as the text.
+If there are no characters in the frame, we make predictions using the same rule as “shortest distance”.
+In these two approaches, the amount of $r$ is the same as that of $\mathtt{text}$.
+Therefore, there is no need to select the top $K$ predictions.
+Appendix D Ablation Studies
+To evaluate our proposed method, we conducted ablation studies and ran our model under the following four patterns.
+  • Rule-based: methods discribed in Section C. In “frame distance”, the frames are obtained from our frame detector.
+  • Only w/ union feature: a deep learning-based simple baseline that calculates relationship scores only using the union feature $\textbf{f}_{i,j}$ obtained in Eq.
+4, without using the fusion module $f_{\mathrm{fuse}}$ in Eq.
+2.
+  • Proposed w/o frame: predict speakers without frame information. In this case, the process is the same as the original SGG process.
+  • Proposed w/ detected frame: use the frame bounding boxes detected by Faster-RCNN. This scenario is a practical use case. We detect both the frame positions and the reading orders.
+  • Proposed w/ given frame: use the frame bounding boxes from Manga109’s annotations. This means that ground truth bounding boxes of frames are available. We then estimate the frame orders.
+Recall@(#text) under three SGG tasks is shown in Table BBD.
+From these tables, we can observe a significant improvement for the model with frame information over the model without it.
+Moreover, the model with given frames performs better than the model with detected frames overall, but “Proposed w/ detected frame” overperforms “Proposed w/ given frame” on Hard.
+It may be because the speaker and the text are not in the same frame in data on Hard, while our frame detector divided frames into smaller ones, making the estimated reading order more accurate.
+Appendix E Experiments with Different Styles of Comics
+To see the behavior of our model on different styles of comics or languages, we tested our model on eBDtheque [20], which comprises 100 images from French, American, and Japanese comics.
+Results under PredCls is shown in Table D.
+From the table, it may be observed that although our model was trained on Manga109, it also performed well on eBDtheque.
+Our method outperformed “shortest distance” but did not perform as well as “frame distance”.
+This was mainly because, in contrast to Japanese manga, the speaker and the text are almost always in the same frame in relatively old European and American comics.
+This experiment highlights the variations in comic styles across different countries and outlines directions for future research.
+Table A: Recall@(#text) for PredCls.
+Manga109Dialog
+Method Easy Hard Total
+Rule-based
+      Short distance 71.43 22.72 63.41
+      Frame distance 81.55 22.06 71.53
+Deep learning-based
+      Only w/ union feature 83.20 28.50 73.99
+      Proposed w/o frame 84.73 29.55 75.48
+      Proposed w/ detected frame 84.77 30.65 75.69
+      Proposed w/ given frame 85.24 29.50 75.89
+Table B: Recall@(#text) for SGCls.
+Manga109Dialog
+Method Easy Hard Total
+Rule-based
+      Short distance 71.06 23.71 63.09
+      Frame distance 80.81 20.61 70.68
+Deep learning-based
+      Only w/ union feature 82.74 28.41 73.59
+      Proposed w/o frame 84.42 29.12 74.92
+      Proposed w/ detected frame 84.51 29.77 75.32
+      Proposed w/ given frame 84.97 29.23 75.62
+Table C: Recall@(#text) for SGDet.
+Manga109Dialog
+Method Easy Hard Total
+Rule-based
+      Short distance 53.11 14.85 46.67
+      Frame distance 64.49 13.67 55.93
+Deep learning-based
+      Only w/ union feature 67.48 21.31 59.71
+      Proposed w/o frame 67.84 21.43 60.04
+      Proposed w/ detected frame 67.92 24.40 60.54
+      Proposed w/ given frame 68.21 22.96 60.60
+Table D: Experimental results on eBDtheque.
+Method eBDtheque
+Rule-based
+      Short distance 77.34
+      Frame distance 80.27
+Deep learning-based
+      Proposed (SGG) w/o frame 75.20
+      Proposed (SGG) w/ frame 77.45
+Appendix F Experiments with Other SGG Models
+This section explains why we chose MOTIFS as the SGG model for our proposed pipeline.
+In addition to MOTIFS, we ran experiments using other SGG models.
+We changed the LSTMs in the fusion module $f_{\mathrm{fuse}}$ to TreeLSTMs of VCTree [21, 10] and Transformer [22, 12], respectively.
+Experimental results are shown in Table E.
+From the table, we can see that the methods using the SGG models all perform better than the rule-based method and DL simple baseline.
+However, there is not a large gap between the results of the different SGG models.
+It is probably because the relationship is only one class in our experiments, and it is difficult for models such as LSTMs to exert their advantages.
+Table E: Results using different SGG models.
+Manga109Dialog
+Method Easy Hard Total
+Frame distance 81.55 22.06 71.53
+DL simple baseline 83.20 28.50 73.99
+Proposed SGG
+      MOTIFS 84.73 29.55 75.48
+      VCTree 83.30 28.85 74.16
+      Transformer 83.76 28.63 74.51
+Appendix G Qualitative Results
+In this section, we show more examples of prediction results.
+Speaker predictions on Easy, Hard and Total are given in Fig.
+I, Fig.
+J, Fig.
+K and Fig.
+L, which show Recall@(#text) in the task of PredCls.
+The green lines represent correct predictions, while the red lines represent incorrect predictions.
+As we can see, our proposed model shows better performance and can deal with more complex cases, especially when there is at least one character in the same frame as the text.
+In this case, if the speaker is not the closest character to the text (like Fig.
+I and the lower part in Fig.
+K) or the speaker is not in the same frame (like the upper part in Fig.
+K), “frame distance” will surely make a wrong prediction, while our model can ease this problem to some degree.
+Fig. I: Predictions on Easy. Courtesy of Yabuno Tenya, Watanabe Tatsuya, Saijo Shinji. The green lines represent correct predictions, while the red lines represent incorrect predictions.
+Fig. J: Predictions on Hard. Courtesy of Yabuno Tenya, Watanabe Tatsuya.
+Fig. K: Predictions on Total. Courtesy of Tashiro Kimu.
+(a) Frame distance.
+(b) Proposed SGG w/o frame.
+(c) Proposed SGG w/ frame.
+Fig. L: Predictions on Total. Courtesy of Kurita Riku.